OpenStudy (anonymous):

expand the expression- ln 12/5

OpenStudy (anonymous):

Note: \[\ln(\frac{x}{y}) = \ln(x) - \ln(y)\] Therefore:\[\ln(12/5) = \ln(12) - \ln(5)\]